koenko:

C=5/9(F−32) The equations above shows how temperature F, measured in degrees Fahrenheit, relates to a temperature C, measured in degrees Celsius. Based on the equation, which of the following must be true? I. A temperature increase of 1 degree Fahrenheit is equivalent to a temperature increase of 5/9 degrees Celsius. II. A temperature increase of 1 degree Celsius is equivalent to a temperature increase of 1.8 degrees Fahrenheit. III. A temperature increase of 59 degree Fahrenheit is equivalent to a temperature increase of 1 degree Celsius. A. I only. B. II only. C. III only. D. I and II only.

AZ:

Hey there

koenko:

hi

AZ:

Your question just got edited. Are the numbers still correct?

AZ:

Let's take a look at case 1 first We increase the temperature in Fahrenheit by 1, will the temperature in Celsius change by 5/9? let's see C=5/9(F−32) Let's first say that we have 32 Fahrenheit what would you get? C = 5/9 (32- 32) C = ?

AZ:

And after that, let's use 33 Fahrenheit C = 5/9 (33-32) C = ?? Is the difference between the first answer and this one 5/9? If so, that makes condition 1 true

AZ:

@koenko

koenko:

oh okay but this is still confessing sorry i is dumm when it comes to this lol

AZ:

Let's go back to when the temperature is 32 F C = 5/9 (32- 32) what is 32 - 32 = ?

koenko:

oh nvm i get it thank u i read it rong

AZ:

What did you get as the answer? I can check for you :)

koenko:

oh no thanks i emailed it to my teach and she said it was indeed correct

AZ:

If you still don't understand, I can explain it to you?